Given y is two times x, z is three times x, and x + y + 3z = 12. What are the values of x, y and z? (1, 2 and 3?)
Answer by ewatrrr(24785)


Hi,
x + y + 3z = 12 y = 2x and z = 3x
x + 2x + 3(3x) = 12
12x = 12
x = 1
You are absolutely correct: x = 1, y = 2 and z = 3
CHECKING the Answer
1 + 2 + 3*3 = 12